HELP NEEDED..!

Identify the slope-intercept form of the equation of a line which contains the points (–2, –1) and (3, 4).

Answer by stanbon(75887)

Identify the slope-intercept form of the equation of a line which contains the points (–2, –1) and (3, 4).
-----------------------------
slope = (4--1)/(3--2) = 5/5 = 1
----
Form: y = mx+b
4 = 1*3+b
b =1
---
Equation:
y = x+1
=========
Cheers,
Stan H.
=========